OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -3 and a5 = -768?

OpenStudy (anonymous):

a5 = -768 and a1 = -3 but a5/a1 = r^4 ( r being the common ratio of the geometric sequence )

OpenStudy (anonymous):

a5/a1 = (-768/-3) =256 = 4^4 or r^4 = 4^4 hence r= +- 4 and a7= a1 * r^6 , here you know a1 and r .......

OpenStudy (anonymous):

now what?!

OpenStudy (anonymous):

put the values of a1 and r to get a7 a7= a1 * r^6

OpenStudy (anonymous):

would you do r^6 to the other side

OpenStudy (anonymous):

when r=4 r^6 =4^6 =4*4*4*4*4*4 =1024 when r =-4 r^6 =(-4)^6 =-4*-4*-4*-4*-4*-4 =1024

OpenStudy (anonymous):

-12,288 -3,072 12,288 3,072

OpenStudy (anonymous):

is that the answerr